OpenStudy (anonymous):

One half of the employees of Acme Co. earn salaries above $18000 annually. One third of the remainder earns salaries between $15000 and $18000. What part of the staff earn below $15000 a) 1/6 b) 2/3 c) 1/2 d) 1/3

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\large \frac{1}{2}\) -----> above $18K

OpenStudy (anonymous):

yeah..

ganeshie8 (ganeshie8):

One third of the remainder earns salaries between $15000 and $18000.

OpenStudy (anonymous):

yup...

ganeshie8 (ganeshie8):

remainder is the remaining HALF employess oly. Out of that HALF, 1/3rd earn salaried between 15K and 18K

OpenStudy (anonymous):

yup..

OpenStudy (anonymous):

so....??

ganeshie8 (ganeshie8):

So, \(\frac{1}{3} \times \frac{1}{2} \) -----> between 15K and 18K

ganeshie8 (ganeshie8):

remaining are the ones whose salary is below 15K

ganeshie8 (ganeshie8):

remaining folks = \(1 - (\frac{1}{2} + \frac{1}{6})\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

that gives us 1/3 but the answer is A :@

ganeshie8 (ganeshie8):

trust ur method, not the answer

OpenStudy (anonymous):

so 1/3 is right? i was getting this answer every which way and i was confused as to why is it wrong :/

ganeshie8 (ganeshie8):

im getting 1/3 too. answer must be wrong

ganeshie8 (ganeshie8):

1/6 is correct if the question was like below :- One half of the employees of Acme Co. earn salaries above $18000 annually. One third of the employees earns salaries between $15000 and $18000. What part of the staff earn below $15000

OpenStudy (anonymous):

yeah so the answer must be wrong ! thanks a lot you cleared the confusion!

ganeshie8 (ganeshie8):

np :) typos in answer keys is not a new thing... they happen all the time !

OpenStudy (anonymous):

yeah true! :)

ganeshie8 (ganeshie8):

Im glad you took time to verify :)

OpenStudy (anonymous):

i have to, otherwise i would have been confused and keep thinking about this ques again and again! :P

ganeshie8 (ganeshie8):

true lol ;)

OpenStudy (anonymous):

:)